# Synchronization-induced violation of thermodynamic uncertainty relations

### v1

@@ I Introduction
In 1665 Huygens observed the synchronization of two pendulum clocks mounted on a common support [1]. Since then, synchronization has emerged as a universal concept in the theory of dynamical systems, with a broad range of applications in fields ranging from science and engineering to social life [2]. More recently, the phenomenon has been investigated and characterized in quantum systems [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] with, however, only a few studies addressing thermodynamic signatures of synchronization  [18, 19]. Just as thermodynamics started in the 1800s spurred by the industrial revolution, in the same way the miniaturization of devices, and in particular the emergence of new quantum technologies, pushes the field of thermodynamics into new applied and fundamental challenges [20, 21, 22, 23, 24, 25, 26]. In the thermodynamics of small (quantum) systems, fluctuations [27, 28, 29] play a prominent role, and thermodynamic uncertainty relations (TURs), derived within the framework of classical stochastic thermodynamics, establish a lower bound to the amount of dissipation needed to reduce relative energy current fluctuations to a given level [30, 32, 33, 31, 34, 35, 36, 37, 38, 39, 40]. This seminal result motivated the quest for possible mechanisms to violate TURs, and consequently reduce the thermodynamic cost of precision. Routes for TUR violations include breaking of time-reversal symmetry (TRS) [42, 43, 41] and quantum coherences [44, 47, 45, 46, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Although it might be intuitive that synchronization, by locking the relative motion of system constituents, can reduce fluctuations, its possible role in TURs violation has not yet been explored.
@@
In this work, we consider two quantum harmonic oscillators (QHOs) coupled to common thermal baths (see Fig. 1 for a schematic drawing of our model). The couplings to one bath are static and induce not only dissipation but also the emergence of correlations between the two otherwise independent oscillators. At strong damping and low temperatures, the oscillators are synchronized, oscillating at a common frequency and in phase opposition. The oscillators are then in contact with a second thermal bath, with periodically driven couplings breaking TRS. In the synchronization regime, the TURs for the local work currents, i.e., the injected or extracted power of each oscillator, can be strongly violated, despite the fact that the TUR for total power is rigorously proven as never violated. In addition, the finite cutoff frequency for the dynamically coupled bath spectral density, which generally implies non-Markovian effects [60, 61, 62], allows local TURs to be violated in the fast driving diabatic regime. In this regime, the violation of TUR is accompanied by the possibility of extracting finite and sizeable power from one oscillator. These results, already present in a isothermal regime, can benefit from the presence of a temperature gradient, especially in the non-linear regime where TUR violation is achieved in a wide parameter region.
@@
II Results
II.1 General setting
Two uncoupled (no direct coupling) quantum harmonic oscillators ($l=A,B$) in contact with two common thermal reservoirs ($\nu=1,2$) are the subsystems constituting the working medium (WM) of the quantum thermal machine under study, as sketched in Fig. 1. The total Hamiltonian is (we set $\hbar=k_{B}=1$)
@@ II.1 General setting
with $H_{l}=\frac{p_{l}^{2}}{2m}+\frac{1}{2}m\omega_{l}^{2}x_{l}^{2}$ the Hamiltonian of the $l$-th QHO (same mass $m$ but different characteristic frequencies $\omega_{A}$ and $\omega_{B}$). The reservoirs $H_{\nu}$ are modelled with the Caldeira-Leggett approach of quantum dissipative systems [63, 64, 65, 66] as a collection of independent harmonic oscillators, while the system-reservoir interaction $H_{{\rm int},\nu}^{(t)}=\sum_{l=A,B}H_{{\rm int},\nu,l}^{(t)}$ is a bilinear coupling $\sim x_{l}\sum_{k=1}^{+\infty}g^{(l)}_{\nu}(t)c_{k,\nu}X_{k,\nu}$ in the subsystem $x_{l}$ and bath $X_{k,\nu}$ position operators, where $c_{k,\nu}$ describe the coupling strengths weighted by the modulating function $g^{(l)}_{\nu}(t)$ (see Appendix A for details). We assume that the couplings with the $\nu=1$ reservoir are weak and oscillate in time [67, 68, 69, 70], with two independent monochromatic drives of the form
$g_{1}^{(A)}(t)=\cos(\Omega t)~{},\quad g_{1}^{(B)}(t)=\cos(\Omega t+\phi)~{},$ (2)
with $\Omega$ the external frequency and $\phi$ a relative phase. The couplings with the $\nu=2$ reservoir, instead, are static, $g_{2}^{(l)}=1$. Furthermore, the couplings with the $\nu=2$ bath are stronger than those with the $\nu=1$ bath.
The properties of the $\nu$-th bath, including possible memory effects and non-Markovian behaviour [71, 72, 73, 74, 75], are governed by the so-called spectral density [64] ${\cal J}_{\nu}(\omega)\equiv\frac{\pi}{2}\sum_{k=1}^{+\infty}\frac{c^{2}_{k,\nu}}{m_{k,\nu}\omega_{k,\nu}}\delta(\omega-\omega_{k,\nu})~{}$, where $m_{k,\nu}$ and $\omega_{k,\nu}$ are the mass and frequency of the $k$-th modes of the $\nu$-th bath.
@@ II.2 Thermodynamic quantities
In the following, we are interested in thermodynamic quantities in the long time limit, when a periodic steady state has been reached. To characterize the working regime and the performance of the quantum thermal machine, we focus on thermodynamic quantities averaged over the period ${\cal T}=2\pi/\Omega$ of the drives in the off-resonant case with $\omega_{B}<\omega_{A}$. Due to the time-dependent drives in Eq. (2), power can be injected or extracted into/from the subsystems $l=A,B$. The average power associated to the $l$-th subsystem is defined as
$P_{l}\equiv\int_{0}^{{\cal T}}\frac{{\rm d}t}{{\cal T}}{\rm Tr}\left[\frac{\partial H_{{\rm int},1,l}^{(t)}(t)}{\partial t}\rho(t_{0})\right]~{},$ (5)
where we have introduced both temporal and quantum averages, $\rho(t_{0})$ is the total density matrix at the initial time $t_{0}\to-\infty$ (see Appendix A), and the time evolution of operators is considered in the Heisenberg picture. The total power is given by $P=\sum_{l}P_{l}$. It is worth noting that power is associated to the temporal variation of the interaction term and, as such, is only due to the dynamical coupling of the WM to the bath $\nu=1$. Furthermore, with our convention positive sign indicates power injection (or current flow toward the WM) and negative sign means power extraction (or current flow out of the WM). The average heat current associated to the $\nu$-th reservoir is given by $J_{\nu}\equiv-\int_{0}^{{\cal T}}\frac{{\rm d}t}{{\cal T}}{\rm Tr}\big{[}\dot{H}_{\nu}(t)\rho(t_{0})\big{]}$, and the balance relation $P+J_{1}+J_{2}=0$ holds true. We also recall that, in accordance with the second law of thermodynamics, the entropy production rate $\dot{S}\equiv-\sum_{\nu}J_{\nu}/T_{\nu}$ is always $\dot{S}\geq 0$ [77, 76].
All quantities undergo fluctuations, and the latter, once averaged over the period of the drives, can be written as [43, 78]
@@
and the associated TUR reads $Q_{O}\geq 2$. Therefore, any mechanism leading to a violation of the TUR results in $Q_{O}<2$. Among all the possible mechanisms for TUR violation we consider the breaking of TRS [42, 43], which in our model is guaranteed by the presence of two independent drives with a finite phase shift $\phi\neq 0$. Notice that breaking of TRS is a necessary but not sufficient condition for TUR violation. Indeed, in the present setup it is possible to prove that the TUR for the total power $P$ is never violated, $Q_{P}\geq 2$ always, regardless of the value of $\phi$ and of the considered spectral densities (see Appendices B and C for a proof of this result). However, as we will show below, this is not the case for the properties of subsystems $l=A,B$, and TUR violation can be achieved by looking at $Q_{P_{l}}$ associated to the subsystem power.
II.3 Synchronization
As stated above, we assume weak (time-dependent) couplings with the $\nu=1$ reservoir. Therefore, the dynamics and all thermodynamic quantities will be evaluated at the lowest order in a perturbative expansion in the system-reservoir interaction $H_{{\rm int},1}^{(t)}$ [68, 70]. Conversely, the static couplings with the reservoir $\nu=2$ are treated at all order in the coupling strength. This is encoded in the response function of the two QHOs that, due to the composite nature of the WM, has a matrix structure. It is now worth to recall that in the resonant case ($\omega_{A}=\omega_{B}$) with static couplings with the $\nu=2$ reservoir, symmetry arguments lead to a dissipation-free subspace (associated to the relative coordinate normal mode $x_{A}-x_{B}$), preventing the system from reaching a stationary regime [70, 79, 80, 81, 82]. For this reason, from now on we consider $\omega_{A}\neq\omega_{B}$ only. Assuming a strictly Ohmic spectral density ${\cal J}_{2}(\omega)=m\gamma_{2}\omega$, the two-by-two response matrix is given by [70]
@@ II.3 Synchronization
The response matrix is a key quantity since it determines the long-time dynamics and it enters into the expressions of all the thermodynamic quantities of interest (see below). In particular, at sufficiently strong damping $\gamma_{2}$ a frequency– and phase–locked mode appears. Indeed, $\gamma_{2}$ not only determines dissipation, but it also mediates correlations between the two, otherwise independent, subsystems. There, the two subsystems $A$ and $B$ become synchronized, oscillating at a common frequency $\bar{\omega}=\sqrt{(\omega_{A}^{2}+\omega_{B}^{2})/2}$ and in phase opposition. The appearance of a common frequency $\bar{\omega}$ can be inferred looking at the eigenvalues of the imaginary part of the response matrix of Eq. (8) ${\bm{\chi}}_{2}(\omega)={\bm{\chi}}_{2}^{\prime}(\omega)+i{\bm{\chi}}_{2}^{\prime\prime}(\omega)$, for different damping strengths $\gamma_{2}$. This is illustrated in Fig. 2(a) where the finite eigenvalue is reported for two different damping strengths. At weak damping $\gamma_{2}\ll\omega_{l}$ two peaks around $\omega_{A}$ and $\omega_{B}$ are present, while at strong damping $\gamma_{2}\gg\omega_{l}$ a unique common frequency at $\bar{\omega}$ is the dominant one. In this regime, the corresponding eigenvecotors show that the two QHOs are in phase opposition.
@@
Mutual synchronization between the two QHOs arises when, regardless of their detuning, they start to oscillate coherently at a common frequency. In the present setup, therefore, synchronization can occur at strong damping $\gamma_{2}\gg\omega_{l}$. To quantify synchronization, we consider the dynamics of local observables and the corresponding Pearson coefficient [83]. Focusing on the position operator of the QHOs, this indicator reads
@@
with $\tilde{\rho}=\exp\{-[\sum_{l=A,B}H_{l}+H_{2}+H_{{\rm int},2,l}]/T_{2}\}$. The Pearson coefficient takes values between $-1$ and $+1$, respectively denoting perfect temporal anti-synchronization and synchronization of the local observables. The value 00, instead, denotes the absence of synchronization [83]. In Fig. 2(b) the behaviour of the Pearson coefficient for our setup with different values of the damping strength $\gamma_{2}$ is reported. This clearly shows that at sufficiently strong damping the two subsystems become synchronized and in anti-phase, reaching ${\cal C}\to-1$ at low temperature. At high temperature this feature is smeared out, with Pearson coefficient ${\cal C}\to 0$. In the opposite, weak damping regime instead ${\cal C}$ is always small, indicating no synchronization.
II.4 TUR violation for local power
In the following, we will investigate the local powers and associated TURs. We will focus on the regime of strong damping and emphasize the role of synchronization to find parameter regions where useful, nonvanishing subsystem power ($P_{l}<0$ and sizeable magnitude) can be obtained with high accuracy ($Q_{P_{l}}\ll 2$).
At the lowest perturbative order in the $\nu=1$ WM-bath interaction, the subsystem power contributions can be written as (see Appendix B for details)
@@ II.4 TUR violation for local power
×$\quad\quad\quad\times\left[{\chi_{2}^{(l,l)}}^{\prime\prime}(\omega)+\cos(\phi){\chi_{2}^{(l,{\bar{l}})}}^{\prime\prime}(\omega)\right],$ (13)
$\delta P_{A/B}\!=\!\mp\Omega\sin(\phi)\int_{-\infty}^{+\infty}\!\!\!\frac{{\rm d}\omega}{4\pi m}\bigg{[}{\cal J}_{1}(\omega+\Omega){\chi_{2}^{(A,B)}}^{\prime}\!(\omega)$
×$\!\!\times\!\coth\!\left(\!\frac{\omega\!+\!\Omega}{2T_{1}}\!\right)\!\!-\!m(\omega\!+\!\Omega)\gamma_{1}^{\prime\prime}\!(\omega\!+\!\Omega){\chi_{2}^{(A,B)}}^{\prime\prime}\!\!(\omega)\!\coth\!\left(\!\frac{\omega}{2T_{2}}\!\right)\!\!\bigg{]}\!,$ (14)
where the last line involves the imaginary part of the damping kernel of Eq. (3) in Fourier space, $\gamma_{1}(\omega)=\gamma_{1}^{\prime}(\omega)+i\gamma_{1}^{\prime\prime}(\omega)$, and we have introduced the function $N(\omega,\Omega)=\coth(\frac{\omega+\Omega}{2T_{1}})-\coth(\frac{\omega}{2T_{2}})$. It is worth to stress that the total power is given by $P=P_{A}^{(0)}+P_{B}^{(0)}$,
$P=-\Omega\int_{-\infty}^{+\infty}\!\frac{{\rm d}\omega}{4\pi m}\mathcal{J}_{1}(\omega+\Omega)N(\omega,\Omega)$
×$\times\!\Big{[}{\chi_{2}^{(A,A)}}^{\prime\prime}\!(\omega)\!+\!{\chi_{2}^{(B,B)}}^{\prime\prime}\!(\omega)\!+\!2\cos(\phi){\chi_{2}^{(A,B)}}^{\prime\prime}\!(\omega)\Big{]}~{}.$ (15)
@@
$J_{1}=\int_{-\infty}^{+\infty}\!\frac{{\rm d}\omega}{4\pi m}(\omega+\Omega)\mathcal{J}_{1}(\omega+\Omega)N(\omega,\Omega)$
×$\times\!\Big{[}{\chi_{2}^{(A,A)}}^{\prime\prime}\!(\omega)\!+\!{\chi_{2}^{(B,B)}}^{\prime\prime}\!(\omega)\!+\!2\cos(\phi){\chi_{2}^{(A,B)}}^{\prime\prime}\!(\omega)\Big{]},$ (16)
and $J_{2}=-P-J_{1}$. It is worth to underline that both $P$ and $J_{\nu}$ have an even dependence on $\phi$ and depend on the imaginary part of the response functions $\chi_{2}^{(l,l^{\prime})}$ only. Importantly, the subsystem power has a new term $\delta P_{l}$ that contains an odd contribution in the phase $\phi$. As we discuss in a moment, this will play a crucial role in determining the TUR violation for the subsystems $l=A,B$ since it represents an explicit TRS-breaking term. Finally, regarding the fluctuations associated to the subsystem power at the lowest perturbative order (see Appendix B) one gets
$D_{P_{l}}=$ $\Omega^{2}\int_{-\infty}^{+\infty}\frac{{\rm d}\omega}{4\pi m}\mathcal{J}_{1}(\omega+\Omega)N(\omega,\Omega)$
@@
Hereafter, we discuss in detail results for the $l=A$ channel and we exploit the $\phi$ phase degree of freedom, setting it to $\phi=\pi/2$ to maximize the TRS-breaking contribution. Note that analogous results can be obtained for $l=B$ letting $\phi=3\pi/2$. As already mentioned, we assume a strictly Ohmic spectral function for the $\nu=2$ reservoir, ${\cal J}_{2}(\omega)=m\gamma_{2}\omega$, while the $\nu=1$ reservoir has an Ohmic spectral density in the Drude-Lorentz form
@@
with a cut-off frequency $\omega_{c}$. Notice that in the strictly Ohmic regime, when $\omega_{c}$ is the highest energy scale, one recovers a memory-less (local in time) response. Finite cut-off values $\omega_{c}$, instead, would in general imply non-Markovian effects [64, 84]. It is worth to note that finite (and small) values of $\omega_{c}$ can be engineered in the context of quantum circuits [85, 84, 86, 87] and have been already inspected in other related dissipative systems [60, 61, 62, 88]. With the spectral density of Eq. (18) the imaginary part of the damping kernel $\gamma_{1}^{\prime\prime}(\omega)$ appearing in Eq. (14) becomes $\gamma_{1}^{\prime\prime}(\omega)={\cal J}_{1}(\omega)/(m\omega_{c})$.
@@
By direct inspection of Eq. (17), in the isothermal regime the fluctuations associated to $l=A$ reduce to $D_{P_{A}}=\Omega\coth(\Omega/(2T))P_{A}^{(0)}$, regardless of the shape of ${\cal J}_{1}(\omega)$, and

### v2

@@ I Introduction
In 1665 Huygens observed the synchronization of two pendulum clocks mounted on a common support [1]. Since then, synchronization has emerged as a universal concept in the theory of dynamical systems, with a broad range of applications in fields ranging from science and engineering to social life [2]. More recently, the phenomenon has been investigated and characterized in quantum systems [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] with, however, only a few studies addressing thermodynamic signatures of synchronization  [18, 19]. Just as thermodynamics started in the 1800s spurred by the industrial revolution, in the same way the miniaturization of devices, and in particular the emergence of new quantum technologies, pushes the field of thermodynamics into new applied and fundamental challenges [20, 21, 22, 23, 24, 25, 26]. In the thermodynamics of small systems, fluctuations [27, 28, 29] play a prominent role, and thermodynamic uncertainty relations (TURs), derived within the framework of classical stochastic thermodynamics, establish a lower bound to the amount of dissipation needed to reduce relative energy current fluctuations to a given level [30, 32, 33, 31, 34, 35, 36, 37, 38, 39]. This seminal result motivated the quest for possible mechanisms to violate TURs, and consequently reduce the thermodynamic cost of precision. Routes for TUR violations include breaking of time-reversal symmetry (TRS) [41, 42, 40] and quantum coherences [43, 46, 44, 45, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. In addition, standard TURs have recently been generalized [42, 59, 60], including the case of time-dependent driving [61, 62, 63]. Although it might be intuitive that synchronization, by locking the relative motion of system constituents, can reduce fluctuations, its possible role in violating TURs has not yet been explored.
@@
In this work, we consider two quantum harmonic oscillators (QHOs) coupled to common thermal baths (see figure 1 for a schematic drawing of our model). The couplings to one bath are static and induce not only dissipation but also the emergence of correlations between the two otherwise independent oscillators. At strong damping and low temperatures, the oscillators are synchronized, oscillating at a common frequency and in phase opposition. The oscillators are then in contact with another thermal bath, with periodically driven couplings breaking TRS. In this work, this simmetry breaking is the first necessary ingredient for TUR violation [30]. We will also show that two other ingredients, namely synchronization and non-Markovianity, are required to achieve TUR violation together with finite output local power. Indeed, we will discuss that a strong violation for the injected or extracted power of each oscillator can occur, despite the fact that the TUR for total power is rigorously proven as never violated. In addition, the finite cutoff frequency for the dynamically coupled bath spectral density, which generally implies non-Markovian effects [64, 65, 66], allows local TURs to be violated in the fast driving diabatic regime. Importantly, in this regime, the violation of TUR is accompanied by the possibility of extracting finite and sizeable power from one oscillator, exploiting synchronization achieved at strong damping. These results, already present in a isothermal regime, can benefit from the presence of a temperature gradient, especially in the non-linear regime where TUR violation is achieved in a wide parameter region. Finally, we show that local TURs can violate also the generalized bound for time-dependent drives [42, 61, 62, 63], even in the diabatic regime with finite output power.
@@
II General setting
II.1 Model
Two uncoupled (no direct coupling) quantum harmonic oscillators ($l=A,B$) in contact with two common thermal reservoirs ($\nu=1,2$) are the subsystems constituting the working medium (WM) of the quantum thermal machine under study, as sketched in figure 1. The total Hamiltonian is (we set $\hbar=k_{B}=1$)
@@ II.1 Model
with $H_{l}=\frac{p_{l}^{2}}{2m}+\frac{1}{2}m\omega_{l}^{2}x_{l}^{2}$ the Hamiltonian of the $l$-th QHO (same mass $m$ but different characteristic frequencies $\omega_{A}$ and $\omega_{B}$). The reservoirs $H_{\nu}$ are modelled with the Caldeira-Leggett approach of quantum dissipative systems [67, 68, 69, 70, 71] as a collection of independent harmonic oscillators, while the system-reservoir interaction $H_{{\rm int},\nu}^{(t)}=\sum_{l=A,B}H_{{\rm int},\nu,l}^{(t)}$ is a bilinear coupling $\sim x_{l}\sum_{k=1}^{+\infty}g^{(l)}_{\nu}(t)c_{k,\nu}X_{k,\nu}$ in the subsystem $x_{l}$ and bath $X_{k,\nu}$ position operators, where $c_{k,\nu}$ describe the coupling strengths weighted by the modulating function $g^{(l)}_{\nu}(t)$ (see Appendix A for details). Notice that with the apex (t) we indicate the parametric time-dependence due to the presence of external drives. We assume that the couplings with the $\nu=1$ reservoir are weak and oscillate in time [72, 73, 74, 75], with two independent monochromatic drives of the form
$g_{1}^{(A)}(t)=f_{A}(\Omega)\cos(\Omega t)~{},\quad g_{1}^{(B)}(t)=f_{B}(\Omega)\cos(\Omega t+\phi)~{},$ (2)
with $\Omega$ the external frequency, $\phi$ a relative phase, and $f_{l}(\Omega)$ relative weights that in general can depend on the external frequency $\Omega$. The couplings with the $\nu=2$ reservoir, instead, are static, $g_{2}^{(l)}=1$. Furthermore, the couplings with the $\nu=2$ bath are stronger than those with the $\nu=1$ bath.
The properties of the $\nu$-th bath, including possible memory effects and non-Markovian behaviour [76, 77, 78, 79], are governed by the so-called spectral density [68] ${\cal J}_{\nu}(\omega)\equiv\frac{\pi}{2}\sum_{k=1}^{+\infty}\frac{c^{2}_{k,\nu}}{m_{k,\nu}\omega_{k,\nu}}\delta(\omega-\omega_{k,\nu})~{}$, where $m_{k,\nu}$ and $\omega_{k,\nu}$ are the mass and frequency of the $k$-th modes of the $\nu$-th bath.
@@ II.2 Thermodynamic quantities
In the following, we are interested in thermodynamic quantities in the long time limit, when a periodic steady state has been reached. To characterize the working regime and the performance of the quantum thermal machine, we focus on thermodynamic quantities averaged over the period ${\cal T}=2\pi/\Omega$ of the drives in the off-resonant case with $\omega_{B}<\omega_{A}$. Due to the time-dependent drives in equation (2), power can be injected or extracted into/from the subsystems $l=A,B$. The average power associated to the $l$-th subsystem is defined as
$P_{l}\equiv\int_{0}^{{\cal T}}\frac{{\rm d}t}{{\cal T}}{\rm Tr}\left[\frac{\partial H_{{\rm int},1,l}^{(t)}}{\partial t}\rho(t)\right]~{},$ (5)
where we have introduced both temporal and quantum averages, $\rho(t)$ is the total density matrix evolved at time $t$ (see Appendix A).
The total power is given by $P=\sum_{l}P_{l}$. It is worth noting that power is associated to the temporal variation of the interaction term and, as such, is only due to the dynamical coupling of the WM to the bath $\nu=1$. Furthermore, with our convention positive sign indicates power injection (or current flow toward the WM) and negative sign means power extraction (or current flow out of the WM). The average heat current associated to the $\nu$-th reservoir is given by $J_{\nu}\equiv-\int_{0}^{{\cal T}}\frac{{\rm d}t}{{\cal T}}{\rm Tr}\big{[}{H}_{\nu}\dot{\rho}(t)\big{]}$, and the balance relation $P+J_{1}+J_{2}=0$ holds true. We also recall that, in accordance with the second law of thermodynamics, the entropy production rate $\dot{S}\equiv-\sum_{\nu}J_{\nu}/T_{\nu}$ is always $\dot{S}\geq 0$ [81, 80].
All quantities undergo fluctuations and the latter, once averaged over the period of the drives, can be written as [42, 82]
@@
and the standard TUR reads $Q_{O}\geq 2$. Therefore, any mechanism leading to a violation of the TUR results in $Q_{O}<2$. Among all the possible mechanisms for TUR violation we consider the breaking of TRS [41, 42], which in our model is guaranteed by the presence of two independent drives with a finite phase shift $\phi\neq 0$. Notice that breaking of TRS is a necessary but not sufficient condition for TUR violation. Indeed, in the present setup it is possible to prove that the TUR for the total power $P$ is never violated, $Q_{P}\geq 2$ always, regardless of the value of $\phi$ and of the considered spectral densities (see Appendices B and C for a proof of this result). However, as we will show below, this is not the case for the properties of subsystems $l=A,B$, and TUR violation can be achieved by looking at $Q_{P_{l}}$ associated to the subsystem power.
III Synchronization
[…]
$\chi^{(l,l^{\prime})}_{2}(t)\equiv im\theta(t)\langle[x_{l}(t),x_{l^{\prime}}(0)]\rangle~{},$ (8)
[…]
@@ III Synchronization
The response matrix is a key quantity since it determines the long-time dynamics and enters into the expressions of all the thermodynamic quantities of interest (see below). In particular, by inspecting the eigenvalue problem posed by ${\bm{\chi}}_{2}(\omega)$ (see above), at sufficiently strong damping $\gamma_{2}$ a frequency– and phase–locked mode appears. Indeed, $\gamma_{2}$ not only determines dissipation, but it also mediates correlations between the two, otherwise independent, subsystems. There, the two subsystems $A$ and $B$ become synchronized, oscillating at a common frequency $\bar{\omega}=\sqrt{(\omega_{A}^{2}+\omega_{B}^{2})/2}$ and in phase opposition. The appearance of a common frequency $\bar{\omega}$ can be inferred looking at the eigenvalues of the imaginary part of the response matrix of equation (9) ${\bm{\chi}}_{2}(\omega)={\bm{\chi}}_{2}^{\prime}(\omega)+i{\bm{\chi}}_{2}^{\prime\prime}(\omega)$, for different damping strengths $\gamma_{2}$. This is illustrated in figure 2(a) where the finite eigenvalue is reported for two different damping strengths. At weak damping $\gamma_{2}\ll\omega_{l}$ two peaks around $\omega_{A}$ and $\omega_{B}$ are present, while at strong damping $\gamma_{2}\gg\omega_{l}$ a unique common frequency at $\bar{\omega}$ is the dominant one. In this regime, the corresponding eigenvectors show that the two QHOs are in phase opposition.
@@
Mutual synchronization between the two QHOs arises when, regardless of their detuning, they start to oscillate coherently at a common frequency. In the present setup, therefore, synchronization can occur at strong damping $\gamma_{2}\gg\omega_{l}$. To quantify synchronization, we consider the dynamics of local observables and the corresponding Pearson coefficient [87]. It is worth to note that since we are interested in the steady state regime, the Pearson coefficient attains its asymptotic value and does not depend on time. Focusing on the position operator of the QHOs, this indicator reads
@@
with $\tilde{\rho}=\exp\{-[\sum_{l=A,B}H_{l}+H_{2}+H_{{\rm int},2,l}]/T_{2}\}$. The Pearson coefficient takes values between $-1$ and $+1$, respectively denoting perfect temporal anti-synchronization and synchronization of the local observables. The value 00, instead, denotes the absence of synchronization [87]. In figure 2(b) the behaviour of the Pearson coefficient for our setup with different values of the damping strength $\gamma_{2}$ is reported. This clearly shows that at sufficiently strong damping the two subsystems become synchronized and in anti-phase, reaching ${\cal C}\to-1$ at low temperature. At high temperature this feature is smeared out, with Pearson coefficient ${\cal C}\to 0$. In the opposite, weak damping regime instead ${\cal C}$ is always small, indicating no synchronization.
IV TUR violation for local power
In the following, we will investigate the local powers and associated TURs [30]. We will focus on the regime of strong damping and emphasize the role of synchronization to find parameter regions where useful, nonvanishing subsystem power ($P_{l}<0$ and sizeable magnitude) can be obtained with high accuracy. We also recall that we consider a perturbative expansion in the ratios between the $\nu=1$ and $\nu=2$ system-bath coupling strengths. At the lowest perturbative order, the subsystem power contributions can be written as (see Appendix B for details)
@@ IV TUR violation for local power
×$\times\left[f_{l}^{2}(\Omega){\chi_{2}^{(l,l)}}^{\prime\prime}(\omega)+\cos(\phi)f_{l}(\Omega)f_{{\bar{l}}}(\Omega){\chi_{2}^{(l,{\bar{l}})}}^{\prime\prime}(\omega)\right],$ (14)
$\delta P_{A/B}\!=\!\mp\Omega\sin(\phi)f_{A}(\Omega)f_{B}(\Omega)$
×$\quad\quad\times\int_{-\infty}^{+\infty}\!\!\!\frac{{\rm d}\omega}{4\pi m}\bigg{[}{\cal J}_{1}(\omega+\Omega){\chi_{2}^{(A,B)}}^{\prime}\!(\omega)\coth\!\left(\!\frac{\omega\!+\!\Omega}{2T_{1}}\!\right)$
$\quad\quad-m(\omega\!+\!\Omega)\gamma_{1}^{\prime\prime}\!(\omega\!+\!\Omega){\chi_{2}^{(A,B)}}^{\prime\prime}\!\!(\omega)\!\coth\!\left(\!\frac{\omega}{2T_{2}}\!\right)\!\!\bigg{]}\!,$ (15)
where the last line involves the imaginary part of the damping kernel of equation (3) in Fourier space, $\gamma_{1}(\omega)=\gamma_{1}^{\prime}(\omega)+i\gamma_{1}^{\prime\prime}(\omega)$, and we have introduced the function $N(\omega,\Omega)=\coth(\frac{\omega+\Omega}{2T_{1}})-\coth(\frac{\omega}{2T_{2}})$. It is worth to stress that the total power is given by $P=P_{A}^{(0)}+P_{B}^{(0)}$,
$P=-\Omega\int_{-\infty}^{+\infty}\!\frac{{\rm d}\omega}{4\pi m}\mathcal{J}_{1}(\omega\!+\!\Omega)N(\omega,\Omega)\Big{[}f_{A}^{2}(\Omega){\chi_{2}^{(A,A)}}^{\prime\prime}\!(\omega)$
$\!+\!f_{B}^{2}(\Omega){\chi_{2}^{(B,B)}}^{\prime\prime}\!(\omega)\!+\!2\cos(\phi)f_{A}(\Omega)f_{B}(\Omega){\chi_{2}^{(A,B)}}^{\prime\prime}\!(\omega)\Big{]}~{}.$ (16)
@@
$J_{1}\!=\!\!\!\int_{-\infty}^{+\infty}\!\!\!\frac{{\rm d}\omega}{4\pi m}(\omega\!+\!\Omega)\mathcal{J}_{1}(\omega\!+\!\Omega)N(\omega,\Omega)\Big{[}f_{A}^{2}(\Omega){\chi_{2}^{(A,A)}}^{\prime\prime}\!(\omega)$
$\!+\!f_{B}^{2}(\Omega){\chi_{2}^{(B,B)}}^{\prime\prime}\!(\omega)\!+\!2\cos(\phi)f_{A}(\Omega)f_{B}(\Omega){\chi_{2}^{(A,B)}}^{\prime\prime}\!(\omega)\Big{]},$ (17)
and $J_{2}=-P-J_{1}$. It is worth to underline that both $P$ and $J_{\nu}$ have an even dependence on $\phi$ and depend on the imaginary part of the response functions $\chi_{2}^{(l,l^{\prime})}$ only. Importantly, the subsystem power contains a term $\delta P_{l}$ which is odd in the phase $\phi$. As we discuss in a moment, this will play a crucial role in determining the TUR violation for the subsystems $l=A,B$ since it represents an explicit TRS-breaking contribution. Indeed, setting $\phi=0$ TRS is not broken, $\delta P_{l}=0$, and no TUR violation occurs (see Appendix C). Finally, regarding the fluctuations associated to the subsystem power to the lowest perturbative order (see Appendix B) one gets
$D_{P_{l}}=$ $\Omega^{2}f_{l}^{2}(\Omega)\int_{-\infty}^{+\infty}\frac{{\rm d}\omega}{4\pi m}\mathcal{J}_{1}(\omega+\Omega)N(\omega,\Omega)$
@@
As already mentioned, we assume a strictly Ohmic spectral function for the $\nu=2$ reservoir, ${\cal J}_{2}(\omega)=m\gamma_{2}\omega$, while the $\nu=1$ reservoir has an Ohmic spectral density in the Drude-Lorentz form
@@
with a cut-off frequency $\omega_{c}$. Notice that in the strictly Ohmic regime, when $\omega_{c}$ is the highest energy scale, one recovers a memory-less (local in time) response. Finite cut-off values $\omega_{c}$, instead, would in general imply non-Markovian effects [68, 88]. It is worth to note that finite (and small) values of $\omega_{c}$ can be engineered in the context of quantum circuits [89, 88, 90, 91] and have been already inspected in other related dissipative systems [64, 65, 66, 92]. With the spectral density of equation (19) the imaginary part of the damping kernel $\gamma_{1}^{\prime\prime}(\omega)$ appearing in equation (15) becomes $\gamma_{1}^{\prime\prime}(\omega)={\cal J}_{1}(\omega)/(m\omega_{c})$.
Hereafter we discuss in detail results for the $l=A$ channel and we exploit the $\phi$ phase degree of freedom, setting it to $\phi=\pi/2$ to maximize the TRS-breaking contribution. Furthermore we choose the relative weights as $f_{A}(\Omega)=1$ and $f_{B}(\Omega)=(\Omega/\omega_{c})^{r}$ with the parameter $r\geq 0$. As we will see below this is a flexible choice that allows to obtain precise (small $Q_{P_{A}}$) but sizeable local power $P_{A}<0$, together with high efficiency $\eta$ at the same time. Note that analogous results can be obtained for $l=B$ letting $\phi=3\pi/2$ and inverting the choice for the functions $f_{l}(\Omega)$.
@@
By direct inspection of equation (18), in the isothermal regime the fluctuations associated to $l=A$ reduce to $D_{P_{A}}=\Omega\coth(\Omega/(2T))P_{A}^{(0)}$, regardless of the shape of ${\cal J}_{1}(\omega)$, and